If the voltage of a Zn-H+ cell is 0.45 V at 25 C, when [Zn^2+] = 2.0 M and PH2 = 1 atm, calculate the concentration of H+.
If the voltage of a Zn-H cell is 0.45 V at 25°C when [Zn?) - 2.0 M and P -1.0 atm, calculate the concentration of H.
Answer #1

AG given: Ecell = 0.45V The cell reaction: Zn() +24+ (aq) → Zn2caq) +H2 (9) Zn(s) → htcaa) tał Ex= -Ered e-(-0.76) = 0.76V (1given (244] = 2.09 PH2 = 1.0ath we know that E=E -0.0592V loga - 3 Where E- Cell potential non standard) ša cell potential (s0.3142 = log (cit conse) 0:31X2 0.0592 10.472973 - log Categoja) 10.472973 Log Catrasme) 10 = 10 102 2.971481293 #98129x10 =